The mayor of a town has proposed a plan for the annexation of an adjoining bridge. A political study took a sample of 1000 voters in the town and found that 73 % of the residents favored annexation. Using the data, a political strategist wants to test the claim that the percentage of residents who favor annexation is more than 70 % . Determine the P-value of the test statistic. Round your answer to four decimal places.
Answer #1

H0: p = 0.70

Ha: p > 0.70

Test statistics

z = ( - p ) / sqrt ( p( 1 - p) / n )

= ( 0.73 - 0.70) / sqrt ( 0.70 ( 1 - 0.70) / 1000 )

= 2.07

p-value = P(Z > z)

= 1 - P(Z < z)

= 1 - P(Z < 2.07)

= 1 - 0.9808 (from Z table )

= 0.0192